# Reinterpreting Landauer conductance, solving the quantum measurement problem, grand unification

### v1

@@ 3 Conclusion
AD of a scattering matrix element like $t_{31}$ or $s_{\alpha\gamma}$ can show two kinds of topological behavior. In one case the AD closes on one side of the complex plane $(Im[t_{31}],Re[t_{31}])$ without enclosing the singularity (0,0). In that case both $\rho_{lpd}$ and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ will oscillate with positive-negative values and will be equal in magnitude. Thus the equality derived in Ref. [8] make sense and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ is determined by $\rho_{lpd}$. However, if the AD of $t_{31}$ does enclose the singularity in the plane $(Im[t_{31}],Re[t_{31}])$ then the equality of the integrations in Eq. 62 itself breaks down because of topological reasons. Then the integrands will still be comparable in magnitudes as derived in Ref. [8] but they will be qualitatively different, not resembling each other. The LHS of Eq. 62 not being a winding number will always be zero while the RHS being a winding number will be $2\pi$. Never the less, $\rho_{lpd}$ still determines $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ as they are both expressions of the same AD, however $\rho_{lpd}$ will oscillate on the positive side only while $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ will oscillate with positive and negative values. One can experimentally determine this AD [10] and by carefully examining the AD one can infer the behavior of $\rho_{lpd}$ and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ and relate them. However, it is easy to tune a mesoscopic system into a regime with only Fano resonances and thereby obtain an equality between $\rho_{lpd}$ and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ for any system. Thus in principle a measurable quantity $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ is determined by a number of local partial states $\rho_{lpd}(\textbf{r})$ inspite of all possible linear superposition of states at the point r is demonstrated and hence solves the quantum measurement problem in the extreme quantum regime as well. Since $\rho_{lpd}(\textbf{r})$ is thereby real, and is related to a local time that is fully consistent with relativity as well as quantum mechanics, hence unifies the two. Prof. Roger Penrose has been repeatedly stating in his public talks that since outcome of a quantum measurement within the axiomatic framework of quantum mechanics is random it will not be possible to unify relativity and quantum mechanics unless one can show that quantum measurements are deterministic. That precisely come out in our work. The $\rho_{lpd}$ cannot be defined in the axiomatic framework of quantum mechanics, but can deterministically give the outcome of a quantum measurement, and naturally solve the problem of unification.

### v2

@@ 3 Conclusion
AD of a scattering matrix element like $t_{31}$ or $s_{\alpha\gamma}$ can show two kinds of topological behavior. In one case the AD closes on one side of the complex plane $(Im[t_{31}],Re[t_{31}])$ without enclosing the singularity (0,0). In that case both $\rho_{lpd}$ and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ will oscillate with positive-negative values and will be equal in magnitude. Thus the equality derived in Ref. [8] make sense and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ is determined by $\rho_{lpd}$. However, if the AD of $t_{31}$ does enclose the singularity in the plane $(Im[t_{31}],Re[t_{31}])$ then the equality of the integrations in Eq. 63 itself breaks down because of topological reasons. Then the integrands will still be comparable in magnitudes as derived in Ref. [8] but they will be qualitatively different, not resembling each other. The LHS of Eq. 63 not being a winding number will always be zero while the RHS being a winding number will be $2\pi$. Never the less, $\rho_{lpd}$ still determines $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ as they are both expressions of the same AD, however $\rho_{lpd}$ will oscillate on the positive side only while $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ will oscillate with positive and negative values. One can experimentally determine this AD [10] and by carefully examining the AD one can infer the behavior of $\rho_{lpd}$ and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ and relate them. However, it is easy to tune a mesoscopic system into a regime with only Fano resonances and thereby obtain an equality between $\rho_{lpd}$ and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ for any system. Thus in principle a measurable quantity $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ is determined by a number of local partial states $\rho_{lpd}(\textbf{r})$ in spite of all possible linear superposition of states at the point r is demonstrated and hence solves the quantum measurement problem in the extreme quantum regime as well. Since $\rho_{lpd}(\textbf{r})$ is thereby real, and is related to a local time that is fully consistent with relativity as well as quantum mechanics, hence unifies the two. Prof. Roger Penrose has been repeatedly stating in his public talks that since outcome of a quantum measurement within the axiomatic framework of quantum mechanics is random it will not be possible to unify relativity and quantum mechanics unless one can show that quantum measurements are deterministic. That is basically because if the theory of quantum mechanics fail to give a mechanism for measurement outcomes at a regime where it works the best, that is the micro world, then it cannot give a measurement outcome in the limit when the macro world evolves from the micro world. Our solution to the problem of quantum measurement and unification of classical and quantum worlds corroborates Penrose’s intuition. The $\rho_{lpd}$ cannot be defined in the axiomatic framework of quantum mechanics, but can deterministically give the outcome of a quantum measurement, and naturally solve the problem of unification.